OpenStudy (anonymous):

Solve. (x + 2)^2 = 1

OpenStudy (anonymous):

(x+2)^(2)=1 Take the square root of each side of the equation to setup the solution for x. ~((x+2)^(2))=\~(1) Remove the perfect root factor (x+2) under the radical to solve for x. (x+2)=\~(1) Pull all perfect square roots out from under the radical. In this case, remove the 1 because it is a perfect square. (x+2)=\1 First, substitute in the + portion of the \ to find the first solution. (x+2)=1 Remove the parentheses around the expression x+2. x+2=1 Since 2 does not contain the variable to solve for, move it to the right-hand side of the equation by subtracting 2 from both sides. x=-2+1 Add 1 to -2 to get -1. x=-1 Next, substitute in the - portion of the \ to find the second solution. (x+2)=-1 Remove the parentheses around the expression x+2. x+2=-1 Since 2 does not contain the variable to solve for, move it to the right-hand side of the equation by subtracting 2 from both sides. x=-2-1 Subtract 1 from -2 to get -3. x=-3 The complete solution is the result of both the + and - portions of the solution. x=-1,-3